OpenStudy (anonymous):

Two less than one-sixth of a certain number is 1 more than one-fourth that number. What is the number?

OpenStudy (anonymous):

\[(1/6)x -2 = (1/4)x+1\] Times both sides by 24 to eliminate fractions: \[4x-48=6x+24\] Solve: \[-72=2x\] \[x=-36\]

OpenStudy (anonymous):

Plugging back into original formula you get -8 on both sides.

OpenStudy (anonymous):

(1/6)x-2=1+(1/4)x 2x-24=12+3x (here I multiplied by 12 to get rid of fractions) so then you isolate x on one side and get -24=12+x then just solve for x x=-36 1/6 of -36 is -6. -6-2 =-8 1/4 of -36 is -9. -9+1 =-8